3. The distance travelled by a train on a straight track in the first two seconds is given by

𝑠= ∫20 20(1βˆ’π‘’^βˆ’π‘‘)𝑑𝑑20

Find the distance travelled in Metres


1
Expert's answer
2021-05-31T12:23:18-0400

"\mathrm{Take\:the\:constant\:out}:\quad \int a\cdot f\left(x\right)dx=a\cdot \int f\left(x\right)dx"

"=20\cdot \int _0^{-2}-1+e^udu"

"\int _a^bf\left(x\right)dx=-\int _b^af\left(x\right)dx,\:a<b"

"=20\left(-\int _{-2}^0-1+e^udu\right)"

"=20\left(-\left(-\int _{-2}^01du+\int _{-2}^0e^udu\right)\right)"

"=20\left(-\left(-2+1-\frac{1}{e^2}\right)\right)"

"=-20\left(-\frac{1}{e^2}-1\right)"

"22.70670"m